OpenStudy (anonymous):

Suppose lightening occurs 1.00mi away (about 1609m) . How much time elapses between seeing the lightening and hearing the thunder it causes? The speed of light is constant in air: v=3.00 x 10^8 m/s.

OpenStudy (anonymous):

Use\[v=\frac{ \Delta d }{ \Delta t }\]You are given delta d and v. Can you rearrange and solve for delta t?

OpenStudy (anonymous):

You will also need to know the speed of sound in air.

OpenStudy (anonymous):

I did 3 x 10^8 = 1609/t and it came out as 5.36 x 10^6 >_< DId I plug in the variables wrong? I know the answer is 4.69 s because my teacher gives us the answers, im just not sure how to get it.

OpenStudy (anonymous):

oh and rearranged I did 3 x 10^8(t) = 1609

OpenStudy (anonymous):

So\[t=\frac{ 1609 }{ 3\times 10^8 }=?\]This will tell you how long it takes the light from the lightning to reach your eyes. But you also need to use the same formula to calculate the time it takes for the sound of the thunder to reach your ears. And for that you'll need the speed of sound in air.

OpenStudy (anonymous):

how do I find that? or should I look it up?

OpenStudy (anonymous):

I assumed that your teacher must have provided that information. The speed of sound in air varies with atmospheric conditions but under standard conditions is about 343 m/s.

OpenStudy (anonymous):

I found 333.4+0.6T_c m/s on google

OpenStudy (anonymous):

no. Everything he provided I posted above

OpenStudy (anonymous):

Correct. If you know the air temperature, this formula will give you v. Most of the time, we assume dry air at 20C, giving 345 m/s.

OpenStudy (anonymous):

Using this value for v, you can calculate how long it takes for the sound to reach your ears.

OpenStudy (anonymous):

got it! thank you :)

OpenStudy (anonymous):

You're welcome.